# Uncertainty-aware Distributional Offline Reinforcement Learning

### v1

@@ 4.4 Training Algorithm
  Initialize critic parameter $\theta$, actor parameter $\phi$, perturbation model $\xi$, quantile huber loss threshold $\mathcal{K}$, number of generated quantiles $N,K$, diffusion parameter $\psi$, distortion $\mathbb{D}$, learning rate $\eta$.
  Sample quantile $\tau_{i},\tau_{j}$, where $i=\{0,\cdots,N\},j=\{0,\cdots,K\}$ from $\mathcal{U}(0,1)$
  for $t=1,\cdots$ do
     Sample $B$ transitions $(s,a,r,s^{\prime})$ from the offline dataset
     Compute the $\delta$ based on Equation 3
     Sample $\beta$ from $d_{\psi}(s,a)$ and compute the policy $\pi_{\phi}$ based on Equation 7
     Compute critic loss by using Equation 5
     Compute the loss for $\xi$ by using Equation 6
     Compute the diffusion loss by using Equation 11 and classifier’s loss
     Compute the loss of actor by adding the loss of $\xi$ and the diffusion model
     Update $\theta$, $\phi$, $\psi$ accordingly
  end for
@@
The critic seeks to learn the return distribution, while the diffusion model aims to comprehend the behavioral policy, serving as a baseline action for the actor to implement risk-averse perturbations. As aforementioned, the distortion operator $\mathbb{D}$ is required. It can be any distorted expectation objective, such as CVaR, Wang, and CPW. In this work, we focus on $\text{CVaR}_{\alpha=0.1}$, which is the majority distorted measure used in recent literature (Rigter et al., 2022; Urpí et al., 2021; Ma et al., 2021).
5 Experiments
@@
5.1 Risk-Sensitive D4RL
@@
Tasks Firstly, we test the performance of the proposed UDAC in risk-sensitive D4RL environments. Hence, we turn our attention to stochastic D4RL (Urpí et al., 2021). The original D4RL benchmark (Fu et al., 2020) comprises datasets obtained from SAC agents that exhibit varying degrees of performance (Medium, Expert, and Mixed) on the Hopper, Walker2d, and HalfCheetah in MuJoCo environments (Todorov et al., 2012). Stochastic D4RL modifies the rewards to reflect the stochastic damage caused to the robot due to unnatural gaits or high velocities; please refer to Section D.1 for details. The Expert dataset consists of rollouts generated by a fixed SAC agent trained to convergence; the Medium dataset is constructed similarly, but the agent is trained to attain only 50% of the expert agent’s return. We have selected several state-of-the-art risk-sensitive offline RL algorithms as well as risk-neutral algorithms. It includes: O-RAAC (Urpí et al., 2021), CODAC (Ma et al., 2021), O-WCPG (Tang et al., 2020), Diffusion-QL (Wang et al., 2023), BEAR (Kumar et al., 2019) and CQL (Kumar et al., 2020). Note that, the O-WCPG is not originally designed for offline settings. We follow the same procedure as described in Urpí et al. (2021) to transform it into an offline algorithm. Among them, ORAAC, CODAC and O-WCPG are risk-sensitive algorithms, Diffusion-QL, BEAR and CQL are risk-natural. Moreover, ORAAC and CODAC are distributional RL algorithms.
@@
Results In Table 1, we present the mean and $\text{CVaR}_{0.1}$ returns on test episodes, along with the duration for each approach. These results are averaged over 5 random seeds. As shown, UDAC outperforms the baselines on most datasets. It is worth noting that CODAC contains two different variants that optimize different objectives. In this task, we select CODAC-C as the baseline since it targets risk-sensitive situations. Furthermore, we observe that the performance of CQL varies significantly across datasets. Under the risk-sensitive setting, it performs better than two other risk-sensitive algorithms, O-RAAC and CODAC, in terms of $\text{CVaR}_{0.1}$. One possible explanation for this observation is that CQL learns the full distribution, which helps to stabilize the training and thus yields a better $\text{CVaR}_{0.1}$ than O-RAAC and CODAC.
@@
Table 1: Performance of the UDAC algorithm on risk-sensitive offline D4RL environments using medium, expert and mixed datasets. We compare the $\text{CVaR}_{0.1}$ and mean of the episode returns, and report the mean (standard deviation) of each metric. Our results indicate that across all environments, UDAC outperforms benchmarks with respect to $\text{CVaR}_{0.1}$. Furthermore, in environments that terminate, UDAC achieves longer episode durations as well.
@@
$\text{CVaR}_{0.1}$ Mean $\text{CVaR}_{0.1}$ Mean $\text{CVaR}_{0.1}$ Mean
@@
5.2 Risky Robot Navigation
Tasks We now aim to test the risk-averse performance of our proposed method. Following the experimental setup detailed in Ma et al. (2021), we will conduct experiments on two challenging environments, namely Risky Ant and Risky PointMass. In these environments, an Ant robot must navigate from a randomly assigned initial state to the goal state as quickly as possible. However, the path between the initial and goal states includes a risky area, which incurs a high cost upon passing through. While a risk-neutral agent may pass through this area, a risk-sensitive agent should avoid it. For further information on the experimental setup, including details on the environments and datasets, refer to D.1. We perform the comparison with the same baselines as the previous section. We present a detailed comparison between CODAC and ORAAC, two state-of-the-art distributional risk-sensitive offline RL algorithms, as well as other risk-neutral algorithms. Our discussion covers a detailed analysis of their respective performance, strengths, and weaknesses.
Results The effectiveness of each approach is evaluated by running 100 test episodes and measuring various metrics, including the mean, median, and $\text{CVaR}_{0.1}$ returns, as well as the total number of violations (i.e., the time steps spent inside the risky region). To obtain reliable estimates, these metrics are averaged over 5 random seeds. Moreover, we also provide a visualization of the evaluation trajectories on the Risky Ant environment on Figure 1, and we can see that the CODAC and ORAAC do have some risk-averse action, while UDAC successfully avoids the majority of the risk area. We also note that the Diffusion-QL failed to avoid the risk area as it is not designed for risk-sensitive tasks.
Table 2: Risky robot navigation quantitative evaluation.
@@ 5.2 Risky Robot Navigation
Mean Median $\text{CVaR}_{0.1}$ Violations Mean Median $\text{CVaR}_{0.1}$ Violations
@@
Performance of UDAC Our experimental findings indicate that UDAC consistently outperforms other approaches on both the $\text{CVaR}_{0.1}$ return and the number of violations, suggesting that UDAC is capable of avoiding risky actions effectively. Furthermore, UDAC exhibits competitive performance in terms of mean return, attributed to its superior performance in $\text{CVaR}_{0.1}$. A noteworthy observation is that in the Risky PointMass environment, UDAC learns a safe policy that avoids the risky region entirely, resulting in zero violations, even in the absence of such behavior in the training dataset. In addition, in Section 5.5, we demonstrate that UDAC is effective in optimizing alternative risk-sensitive objectives such as Wang and CPW. Our results illustrate that UDAC can achieve comparable performance to UDAC on $\text{CVaR}_{0.1}$ return and the number of violations in the presence of distorted rewards.
@@
Comparison to ORAAC ORAAC and UDAC both optimize the CVaR objective, but ORAAC employs imitation learning to constrain the learned policy to the empirical behavior policy. However, this approach may result in suboptimal performance due to the presence of many suboptimal trajectories in the dataset and the risk-neutral nature of the dataset. This issue is particularly relevant in offline RL, where the goal is to leverage large datasets for training that are often heterogeneous in terms of the quality of the behavior policy. In contrast, UDAC is better suited to learn from such datasets by using the diffusion approach, as demonstrated in our results, which show that it outperforms ORAAC in these settings.
Comparison to CODAC CODAC employs a strategy that is similar to ORACC by introducing a penalty constant to discourage out-of-distribution actions. Additionally, CODAC adopts a pre-defined behavior policy to avoid using low-quality behavioral policies in the datasets. However, this approach may lead to suboptimal performance due to the absence of some edge cases. In contrast, UDAC aims to learn to model the behavior policy from the datasets directly via denoising, which has better generalizability. The results suggest that the UDAC can avoid most of the risky areas.
Comparsion to Diffusion-QL In environments with significant risks, Diffusion-QL tends to learn policies that exhibit poor tail-performance, which is evident from its low $\text{CVaR}_{0.1}$ performance despite the high median performance. Additionally, the mean performance of Diffusion-QL is not satisfactory since it can be heavily influenced by outliers that are not represented in the training dataset. On the other hand, in the Risky Ant environment, Diffusion-QL performs poorly on all metrics but achieves a higher median value than ORAAC. Specifically, Diffusion-QL fails to reach the goal and avoid risky behavior. The majority reason is that the behavior policy used in these environments is risk-neutral and Diffusion-QL shows a strong capability to learn the risk-neutral policy.
5.3 Risk-Neutral D4RL
Tasks Next, we demonstrate the effectiveness of UDAC even in scenarios where the objective is to optimize the standard expected return. For this purpose, we evaluate UDAC on the widely used D4RL Mujoco benchmark and compare its performance against state-of-the-art algorithms previously benchmarked in Wang et al. (2023). We include ORAAC and CODAC as offline distributional RL baselines. In our evaluation, we report the performance of each algorithm in terms of various metrics, including the mean and median returns, as well as the $\text{CVaR}_{0.1}$ returns.
Table 3: Normalized Return on the risk-neutral D4RL Mujoco Suite.
@@ 5.3 Risk-Neutral D4RL
Results As per Wang et al. (2023), we directly report results for non-distributional approaches. To evaluate CODAC and ORAAC, we follow the same procedure as Wang et al. (2023), training for 1000 epochs (2000 for Gym tasks) with a batch size of 256 and 1000 gradient steps per epoch. We report results averaged over 5 random seeds, as presented in Table 3, demonstrate that UDAC achieves remarkable performance across all 9 datasets, outperforming the state-of-the-art on 5 datasets (halfcheetah-medium, halfcheetah-medium-replay, walker2d-medium-replay, hopper-medium-expert, and walker2d-medium-expert). Although UDAC performs similarly to Diffusion-QL in risk-neutral tasks, it is designed primarily for risk-sensitive scenarios, and its main strength lies in such situations. In contrast, UDAC leverages diffusion inside the actor to enable expressive behavior modeling and enhance performance compared to CODAC and ORAAC.
5.4 Hyperparameter Study in Risk-Sensitive D4RL
@@
We intend to investigate a crucial hyperparameter within the actor network, namely $\lambda$, in our ongoing research. In Figure 6, we present an ablation study on the impact of this hyperparameter in risk-sensitive D4RL. Our findings indicate that selecting an appropriate value of $\lambda$ is critical to achieving high performance as it balances pure imitation from the behavior policy with pure reinforcement learning. Specifically, as $\lambda\rightarrow 0$, the policy becomes more biased towards imitation and yields poor risk-averse performance, while as $\lambda\rightarrow 1$, the policy suffers from bootstrapping errors leading to lower performance. Our experiments suggest that values of $\lambda$ in the range of $[0.05,0.5]$ tend to perform the best, although the optimal value may vary depending on the specific environment. In all MuJoCo experiments, we set the $\lambda$ parameter, which modulates the action perturbation level, to 0.25 except for the HalfCheetah-medium experiment, where it was set to 0.5. However, as illustrated in Figure 6, our results indicate that these values are not the best for all environments, but instead perform well across most. Due to the page limit, the remaining hyperparameter studies on Risky robot navigation can be found in Section D.2.
5.5 Different Distorted Operators
In order to demonstrate the performance of UDAC in different distorted operators, we have selected the risky environment - Risky PointMass and Risky Ant to conduct the experiments. We conduct the experiments using a similar setting as mentioned in previous parts. And the results can be found on Section D.3. We can find that, the UDAC-Wang achieves similar performance with UDAC in the Risky PointMass but slightly worse on Risky Ant. Based on the conclusions of previous studies (Ma et al., 2021; Dabney et al., 2018a), it can be inferred that Wang’s risk preferences are marginally more inclined towards risk-seeking when compared to the CVaR approach. This inference can be attributed to the fact that Wang assigns a non-zero weight (although negligible) to quantile values that exceed the risk cutoff threshold. On the other hand, the CPW approach can be regarded as being similar to a risk-neutral strategy. This is because the CPW approach is designed to model human game-play behavior, and thus does not explicitly incorporate risk preferences into its framework.
6 Conclusion and Future Work

### v2

@@ IV-C Training Algorithm
1 Initialize critic parameter $\theta$, actor parameter $\phi$, perturbation model $\xi$, quantile huber loss threshold $\mathcal{K}$, number of generated quantiles $N,K$, diffusion parameter $\psi$, distortion $\mathbb{D}$, learning rate $\eta$;
2 Sample quantile $\tau_{i},\tau_{j}$, where $i=\{0,\cdots,N\},j=\{0,\cdots,K\}$ from $\mathcal{U}(0,1)$;
3 for $t=1,\cdots$ do
4       Sample $B$ transitions $(s,a,r,s^{\prime})$ from the offline dataset;
5       Compute the $\delta$ based on Equation 3;
6       Sample $\beta$ from $d_{\psi}(s,a)$ and compute the policy $\pi_{\phi}$ based on Equation 7;
7       Compute critic loss by using Equation 5;
8       Compute the loss for $\xi$ by using Equation 6;
9       Compute the diffusion loss by using Equation 11 $\triangleright$ add the classifier loss we want to control the process ;
10       Compute the loss of actor by adding the loss of $\xi$ and the diffusion model;
11       Update $\theta$, $\phi$, $\psi$ accordingly;
12
13 end for
@@ 11       Update $\theta$, $\phi$, $\psi$ accordingly;
The critic seeks to learn the return distribution, while the diffusion model aims to comprehend the behavioral policy, serving as a baseline action for the actor to implement risk-averse perturbations. As aforementioned, the distortion operator $\mathbb{D}$ is required. It can be any distorted expectation objective, such as CVaR, Wang, and CPW. In this work, we focus on $\text{CVaR}_{\alpha=0.1}$, which is the majority distorted measure used in recent literature [24, 2, 4].
V Experiments
@@
V-A Implementation Details
Actor
As mentioned in the main text, the architecture of the actor is,
$\pi_{\phi}(s)=\lambda\xi(\cdot|s,\beta)+\beta,\beta\sim\pi_{b},$
If we want to integrate the proposed UDAC into an online setting, we can effortlessly remove the $\beta$ and set $\lambda=1$. For our experiments, we construct the behavioral policy as an MLP-based conditional diffusion model, drawing inspiration from [21, 25]. We employ a 3-layer MLP with Mish activations and 256 hidden units to model $\epsilon_{\psi}$. The input of $\epsilon_{\psi}$ is a concatenation of the last step’s action, the current state vector, and the sinusoidal positional embedding of timestep $i$. The output of $\epsilon_{\psi}$ is the predicted residual at diffusion timestep $i$. Moreover, we build two Q-networks with the same MLP setting as our diffusion policy, with 3-layer MLPs with Mish activations and 256 hidden units for all networks. Moreover, as mentioned in the main body, we set $N$ to be small to relieve the bottleneck caused by the diffusion. In practice, we use the noise schedule obtained under the variance preserving SDE [23].
Critic
For the critic architecture, it is built upon IQN [12] but within continuous action space. We are employing a three-layer MLP with 64 hidden units.
Classifier
For the classifier which is used to control the diffusion process which is just a simple three-layer neural network with 256 hidden units. It was also pre-trained on the corresponding datasets.
V-B Experiments Setup
The stochastic MuJoCo environments are set as follows with a stochastic reward modification. It was proposed by [2], and the following description is adapted:
The risky robot navigation is proposed by [4]. The experiments are conducted on a server with two Intel Xeon CPU E5-2697 v2 CPUs with 6 NVIDIA TITAN X Pascal GPUs, 2 NVIDIA TITAN RTX, and 768 GB memory.
V-C Risk-Sensitive D4RL
@@
Firstly, we test the performance of the proposed UDAC in risk-sensitive D4RL environments. Hence, we turn our attention to stochastic D4RL [2]. The original D4RL benchmark [26] comprises datasets obtained from SAC agents that exhibit varying degrees of performance (Medium, Expert, and Mixed) on the Hopper, Walker2d, and HalfCheetah in MuJoCo environments [27]. Stochastic D4RL modifies the rewards to reflect the stochastic damage caused to the robot due to unnatural gaits or high velocities; please refer to Section V-B for details. The Expert dataset consists of rollouts generated by a fixed SAC agent trained to convergence; the Medium dataset is constructed similarly, but the agent is trained to attain only 50% of the expert agent’s return. We have selected several state-of-the-art risk-sensitive offline RL algorithms as well as risk-neutral algorithms. It includes: O-RAAC [2], CODAC [4], O-WCPG [28], Diffusion-QL [6], BEAR [29] and CQL [30]. Note that, the O-WCPG is not originally designed for offline settings. We follow the same procedure as described in [2] to transform it into an offline algorithm. Among them, ORAAC, CODAC and O-WCPG are risk-sensitive algorithms, Diffusion-QL, BEAR and CQL are risk-natural. Moreover, ORAAC and CODAC are distributional RL algorithms.
@@
In Table I, we present the mean and $\text{CVaR}_{0.1}$ returns on test episodes, along with the duration for each approach. These results are averaged over 5 random seeds. As shown, UDAC outperforms the baselines on most datasets. It is worth noting that CODAC contains two different variants that optimize different objectives. In this task, we select CODAC-C as the baseline since it targets risk-sensitive situations. Furthermore, we observe that the performance of CQL varies significantly across datasets. Under the risk-sensitive setting, it performs better than two other risk-sensitive algorithms, O-RAAC and CODAC, in terms of $\text{CVaR}_{0.1}$. One possible explanation for this observation is that CQL learns the full distribution, which helps to stabilize the training and thus yields a better $\text{CVaR}_{0.1}$ than O-RAAC and CODAC.
@@
TABLE I: Performance of the UDAC algorithm on risk-sensitive offline D4RL environments using medium, expert and mixed datasets. We compare the $\text{CVaR}_{0.1}$ and mean of the episode returns, and report the mean (standard deviation) of each metric. Our results indicate that across all environments, UDAC outperforms benchmarks with respect to $\text{CVaR}_{0.1}$. Furthermore, in environments that terminate, UDAC achieves longer episode durations as well.
@@
$\text{CVaR}_{0.1}\uparrow$ Mean $\uparrow$ $\text{CVaR}_{0.1}\uparrow$ Mean $\uparrow$ $\text{CVaR}_{0.1}\uparrow$ Mean $\uparrow$
@@
V-D Risky Robot Navigation
We now aim to test the risk-averse performance of our proposed method. Following the experimental setup detailed in [4], we will conduct experiments on two challenging environments, namely Risky Ant and Risky PointMass. In these environments, an Ant robot must navigate from a randomly assigned initial state to the goal state as quickly as possible. However, the path between the initial and goal states includes a risky area, which incurs a high cost upon passing through. While a risk-neutral agent may pass through this area, a risk-sensitive agent should avoid it. For further information on the experimental setup, including details on the environments and datasets, refer to V-B. We perform the comparison with the same baselines as the previous section. We present a detailed comparison between CODAC and ORAAC, two state-of-the-art distributional risk-sensitive offline RL algorithms, as well as other risk-neutral algorithms. Our discussion covers a detailed analysis of their respective performance, strengths, and weaknesses.
The effectiveness of each approach is evaluated by running 100 test episodes and measuring various metrics, including the mean, median, and $\text{CVaR}_{0.1}$ returns, as well as the total number of violations (i.e., the time steps spent inside the risky region). To obtain reliable estimates, these metrics are averaged over 5 random seeds. Moreover, we also provide a visualization of the evaluation trajectories on the Risky Ant environment on Figure 1, and we can see that the CODAC and ORAAC do have some risk-averse action, while UDAC successfully avoids the majority of the risk area. We also note that the Diffusion-QL failed to avoid the risk area as it is not designed for risk-sensitive tasks.
TABLE II: Risky robot navigation quantitative evaluation.
@@ V-D Risky Robot Navigation
Mean $\uparrow$ Median $\uparrow$ $\text{CVaR}_{0.1}\uparrow$ Violations $\downarrow$ Mean $\uparrow$ Median $\uparrow$ $\text{CVaR}_{0.1}\uparrow$ Violations $\downarrow$
@@
Our experimental findings indicate that UDAC consistently outperforms other approaches on both the $\text{CVaR}_{0.1}$ return and the number of violations, suggesting that UDAC is capable of avoiding risky actions effectively. Furthermore, UDAC exhibits competitive performance in terms of mean return, attributed to its superior performance in $\text{CVaR}_{0.1}$. A noteworthy observation is that in the Risky PointMass environment, UDAC learns a safe policy that avoids the risky region entirely, resulting in zero violations, even in the absence of such behavior in the training dataset.
@@
V-E Risk-Neutral D4RL
Next, we demonstrate the effectiveness of UDAC even in scenarios where the objective is to optimize the standard expected return. For this purpose, we evaluate UDAC on the widely used D4RL Mujoco benchmark and compare its performance against state-of-the-art algorithms previously benchmarked in [6]. We include ORAAC and CODAC as offline distributional RL baselines. In our evaluation, we report the performance of each algorithm in terms of various metrics, including the mean and median returns, as well as the $\text{CVaR}_{0.1}$ returns.
TABLE III: Normalized Return on the risk-neutral D4RL Mujoco Suite.
@@ V-E Risk-Neutral D4RL
As per [6], we directly report results for non-distributional approaches. To evaluate CODAC and ORAAC, we follow the same procedure as [6], training for 1000 epochs (2000 for Gym tasks) with a batch size of 256 and 1000 gradient steps per epoch. We report results averaged over 5 random seeds, as presented in Table III, demonstrate that UDAC achieves remarkable performance across all 9 datasets, outperforming the state-of-the-art on 5 datasets (halfcheetah-medium, halfcheetah-medium-replay, walker2d-medium-replay, hopper-medium-expert, and walker2d-medium-expert). Although UDAC performs similarly to Diffusion-QL in risk-neutral tasks, it is designed primarily for risk-sensitive scenarios, and its main strength lies in such situations. In contrast, UDAC leverages diffusion inside the actor to enable expressive behavior modeling and enhance performance compared to CODAC and ORAAC.
V-F Hyperparameter Study in Risk-Sensitive D4RL
@@
We intend to investigate a crucial hyperparameter within the actor network, namely $\lambda$, in our ongoing research. In Figure 2, we present an ablation study on the impact of this hyperparameter in risk-sensitive D4RL. Our findings indicate that selecting an appropriate value of $\lambda$ is critical to achieving high performance as it balances pure imitation from the behavior policy with pure reinforcement learning. Specifically, as $\lambda\rightarrow 0$, the policy becomes more biased towards imitation and yields poor risk-averse performance, while as $\lambda\rightarrow 1$, the policy suffers from bootstrapping errors leading to lower performance.
VI Conclusion and Future Work